OpenStudy (anonymous):

A soccer team wins 65% of its matches, and 15% of its matches end in a draw. If the team is scheduled to play 20 matches, about how many matches is it expected to lose?

OpenStudy (anonymous):

65% of 20 = 20(0.65) 15% of 20 = 20(0.15)

OpenStudy (anonymous):

100 - 65 = percentage lost. percentage lost(20) = games expected to lose

OpenStudy (anonymous):

thank youuuu!!!!